OpenStudy (princessaurora):

**Will Fan & Medal** The mean of a population being sampled is 64, and and the standard deviation is 6. If the sample size is 50, the standard error of the mean is? . (Round off your answer to the nearest hundredth.)

OpenStudy (luigi0210):

Would it be something like this? \(\Large \sigma _M = \frac{\sigma}{\sqrt{N}} \)?

OpenStudy (princessaurora):

Yeah but what should I plug in?

OpenStudy (luigi0210):

\(\Large \sigma = Standard~ Deviation \) \(\Large N = Sample ~Size \)

OpenStudy (anonymous):

Maleficent~Beastie

OpenStudy (princessaurora):

what is M?

OpenStudy (luigi0210):

\(\LARGE \sigma_M \) is just representing what you're looking for, it's the Standard Error of the Mean.

OpenStudy (princessaurora):

so i dont plug in anything? I'm confused now

OpenStudy (luigi0210):

Nope, it's just as if you y= (something something) You don't do anything to it, it's just symbolizing what you're looking for

OpenStudy (princessaurora):

Ok I see, but how do I do this problem now?

OpenStudy (luigi0210):

Just plug in vales: \(\Large \sigma = 6\) \(\Large N = 50 \) \(\Large \sigma_M = \frac{6}{\sqrt{50}}\)

OpenStudy (luigi0210):

*values

OpenStudy (princessaurora):

\[M=\frac{ 3\sqrt{2} }{ 5s } ?\]

OpenStudy (princessaurora):

then do i solve for s or simplify it?

OpenStudy (luigi0210):

o.O What are you doing? You just need to put \(\Large \frac{6}{\sqrt{50}} \) into a calculator.

OpenStudy (princessaurora):

is it going to be in decimal form?

OpenStudy (princessaurora):

for fraction i got : \[\frac{ 3\sqrt{2} }{ 5 }\]

OpenStudy (princessaurora):

decimal is 0.85

OpenStudy (luigi0210):

Yup, that's all you had to do (:

OpenStudy (princessaurora):

ok thanks!

OpenStudy (luigi0210):

No prob \(\Large \color{green}{\star} \)